On January 1, 2018, the Haskins Company adopted the dollar-value LIFO method for its one inventory pool. The pool’s value on this date was $840,000. The 2018 and 2019 ending inventory valued at year-end costs were $884,000 and $954,000, respectively. The appropriate cost indexes are 1.04 for 2018 and 1.06 for 2019.

Required:
Complete the below table to calculate the inventory value at the end of 2018 and 2019 using the dollar-value LIFO method. (Round "Year end cost index" to 2 decimal places. Round other final answer values to the nearest whole dollars.)

Inventory Layers Converted to Base Year Cost Inventory Layers Converted to Cost Ending Inventory DVL Cost
Date Inventory at Year-End Cost Year-End Cost Index Inventory Layers at Base Year Cost Inventory Layers at Base Year Cost Year-End Cost Index = Inventory Layers Converted to Cost
01/01/2018 = Base = $0
12/31/2018 = Base =
2018 = $0
12/31/2019 = Base =
2018 =
2019 = $0
Answer #1
Inventory layers converted to base year cost Inventory Layers Converted to Cost Ending inventory DVL cost
Date Inventory at year end cost / Year end cost index = Inventory layers at base year cost Inventory layers at base year cost * Year end cost index = Inventory layers converted to cost
01/01/2018 840000 / 1 = 840000 Base 840000 * 1 = 840000 840000
12/31/2018 884000 / 1.04 = 850000 Base 840000 * 1 = 840000
2018 10000 * 1.04 = 10400 850400
12/31/2019 954000 / 1.06 = 900000 Base 840000 * 1 = 840000
2018 10000 * 1.04 = 10400
2019 50000 * 1.06 = 53000 903400